OpenStudy (anonymous):

Circle J has center J(4, -3) and radius 5. What is the measure, in degrees, of the arc with the endpoints A(9, -3) and B(4,2)?

OpenStudy (phi):

do you know about vectors ? if so I would find V= A-J and W= B-J we get V= (9,-3) - (4,-3) = (5,0) and W= (4,2) - (4, -3) = (0,5) we can think of this as translating (sliding) the circle so its center is at (0,0) in other words, we get this picture  you can see the angle must be 90 degrees